# Unraveling the dynamical behaviors in a quasiperiodic mosaic lattice

### v1

@@ Abstract
Quasiperiodic mosaic systems have attracted significant attention due to their unique spectral properties with exactly known mobility edges, which do not vanish even in the large quasiperiodic potential strength region, although the width of energy window of extended states becomes very narrow and decreases with the increase of strength of the quasiperiodic potential. In this work we study the dynamics of a quasiperiodic mosaic lattice and unravel its peculiar dynamical properties. By scrutinizing the expansion dynamics of wave packet and the evolution of density distribution, we unveil that the long-time density distribution display obviously different behaviors at odd and even sites in the large quasiperiodic potential strength region. Particularly, the time scale of dynamics exhibits an inverse relationship with the quasiperiodic potential strength. To understand these behaviors, we derive an effective Hamiltonian in the large quasiperiodic potential strength region, which is composed of decoupled Hamiltonians defined on the odd and even sites, respectively. While all eigenstates of the effective Hamiltonian defined on even sites are localized, the eigenstates of effective Hamiltonian defined on odd sites include both localized and extended eigenstates. Our results demonstrate that the effective Hamiltonian can describe the dynamical behaviors well in the large quasiperiodic potential strength region and provides an intuitive framework for understanding the peculiar dynamical behaviors in the quasiperiodic mosaic lattice.
@@ I Introduction
Quasiperiodic systems, as pivotal platforms for investigating Anderson localization Anderson , Anderson50 , Mirlin , Lee , 2D and mobility edges MEAnderson1 , MEAnderson2 , Mott1 , have attracted considerable attention and extensive exploration. Compared to the Anderson model in which the extended-localized transition and mobility edge only exist in three dimension 2D , MEAnderson1 , MEAnderson2 , Mott1 , Experiment1 , Experiment4 , Experiment5 , quasiperiodic systems can host the the extended-localized transition and mobility edge in one dimension AAmodel , AAHmodel , Cai , MobilityEdge1 , self-consisitent , Experiment2 , Experiment3 , XCXie , XCXie1 , SarmaMobilityEdge1 , SarmaMobilityEdge2 , GaoXL . While the Aubry-Audré (AA) model hosts extended-localized transition AAmodel , AAHmodel at the self-dual point, generalizations of the AA model MobilityEdge1 , self-consisitent , XCXie , XCXie1 , SarmaMobilityEdge1 , SarmaMobilityEdge2 , Deng , longrange , MosaicAA , globaltheory1 , hidden-duality , RG , Critical , DDVu1 , LiuXJ , WangLi or periodically driving the quasiperiodic systems Qin , Driving , Driving1 , Driving2 , Driving3 , Driving4 can give rise to fruitful phenomena of mobility edges. Recently, a class of quasiperiodic mosaic models MosaicAA , in which the quasiperiodic potential periodically occurs with a fixed interval, was proposed as a new type of quasiperiodic systems with exact mobility edges. Different from previous models SarmaMobilityEdge1 , SarmaMobilityEdge2 , the quasiperiodic mosaic models have mobility edges even in the large quasiperiodic potential strength region and can host multiple mobility edges. Stimulated by this quasiperiodic mosaic model, many other mosaic systems have been studied and show quite interesting behaviors LiuYX , Longhi , MosaicSystem1 , MosaicSystem2 , MosaicSystem3 , MosaicSystem4 , MosaicSystem5 , Mosaic6 , Dwiputra . Recently, the experimental realization of the quasiperiodic mosaic model was also reported mosaicAAExperiment .
@@
While the diffusion of a wave packet in a uniform lattice is ballistic, the Anderson localization can suppress the diffusion of the wave packet. For a one-dimensional quasiperiodic lattice with mobility edge, for example, the Ganeshan-Pixley-Das Sarma (GPD) model SarmaMobilityEdge2 , there exists an intermediate regime in which the extended and localized states coexist and are separated by the moblity edges. The dynamical behaviors in this region are quite subtle and have drawn considerable attention, showcasing novel phenomena Dynamics1 , Dynamics2 , Dynamics3 . It was shown that the dynamical behaviors of wave packet in the intermediate regime of the GPD model exhibit a blend of both localized and extended features Dynamics1 , distinguishing them from dynamics in the multifractal region. For quasiperiodic mosaic systems, despite Avila’s global theory providing an exact expression for the mobility edge, the dynamics of the quasiperiodic mosaic model has not yet been studied and is worth further exploration.
@@
In this work, we investigate the dynamics of a quasiperiodic mosaic model across a range of quasiperiodic potential strengths and particularly unravel the peculiar dynamical behavior in the large quasiperiodic potential strength limit. Our analysis involves visually tracking the evolution of a Gaussian wave packet and study the evolution of particle number distribution with different initial states. We observe quite distinct behaviors under different quasiperiodic potential strengths. Particularly, in the large quasiperiodic potential strength limit, we observe significant density distribution differences between even and odd sites, with the expansion timescale of the wave packet showing a linear relationship with the quasiperiodic potential strength $\lambda$. To understand the physical original of these peculiar dynamical behaviors, we derive an effective Hamiltonian in the large quasiperiodic potential strength region, indicating that the effective Hamiltonian for the odd and even sites are decoupled. This effective Hamiltonian provides a clear physical picture of the even-odd difference of the dynamical behaviors and effectively captures the scaling behavior of eigenvalues, as well as the property of extended-localized transition. Additionally, analysis of the eigenstates of the effective Hamiltonian yields further insights into the understanding of the spectral structure of the quasiperiodic mosaic model. Our study sheds light on the physical properties of the quasiperiodic mosaic model, unraveling the physical origin of the peculiar dynamical behaviors in the quasiperiodic mosaic model.
@@
The rest of the paper is organized as follows: In Sec.II, we introduce the quasiperiodic mosaic model and its main properties briefly. Then, we study the dynamics of the wave packet expansion and the evolution of particle number distribution with different initial states. In Sec.III, we derive the effective Hamiltonian in the large quasiperiodic potential strength limit and analyze the properties of effective Hamiltonian, through which we can understand clearly the even-odd difference of the wave packet dynamics. Finally, in Sec. IV, we give a summary of our findings.
@@ II.1 Model
where $\hat{c}_{j}^{\dagger}(\hat{c}_{j})$ is the fermion creation(annihilation) operator and $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the particle number operator at the $j$-th site. $\lambda$ is the strength of quasiperiodic potential and $J$ is the hopping amplitude, for which we set $J=1$ as the unit of energy in the following calculation. $\omega$ is an irrational number and we choose $\omega=\frac{\sqrt{5}-1}{2}$. $\phi$ is a random phase which leads to the realization of different quasiperiodic potentials. We note that the quasiperiodic potential is distributed only on even sites.
@@
where $\psi_{n,j}$ represents the $n$-th eigenstate’s amplitude on the $j$-th site. The fractal dimension $D_{2}$ takes different values at thermodynamic limit in various regions: while $D_{2}\rightarrow 1$ in the extended region, $D_{2}\rightarrow 0$ in the localized region.
@@
Figure 1: Basic properties and wave packet dynamics of the quasiperiodic mosaic model. (a) The eigenvalue spectrum versus $\lambda$ with the fractal dimension of corresponding eigenstate denoted by $D_{2}$. (b) The proportion of extended states in quasiperiodic mosaic model. (c-e) The evolution of wave packet under different quasiperiodic potential strength. From Left to right, we choose $\lambda=0.1,1,10$. From top to bottom, we choose $t=0,10,10^{2},10^{3},10^{4}$. The blue (red) points represent the even (odd) wavefunction distribution. We choose the Gaussian wave packet with $\sigma=1$ as the initial state. In sub-figure (e), the lines represent the even (odd) wavefunction distribution from the evolution of the effective Hamiltonian. We choose $L=987$ in our calculation and take 100 samples.
@@ II.2 Wave packet dynamics
$|\psi(0)\rangle=C\sum_{j}\exp(-(j-j_{c})^{2}/2\sigma^{2})|j\rangle,$
@@
To get an intuitive picture of the evolution of wave packets for various $\lambda$, we show the distribution of particle number at different times $N_{j}(t)=\langle\psi(t)|\hat{n}_{j}|\psi(t)\rangle$. The results are shown in Fig.1(c-e), which exhibit the wave packet dynamics across varying degrees of quasiperiodic potential strength. In Fig. 1 (c), with $\lambda=0.1$, all the eigenstates are extended. The wave packet expands and lose all the initial information after a long time evolution. Here we adopted logarithmic coordinates for $N_{j}(t)$. Notably, there are no discernible differences between distributions at even and odd sites. In Fig.1 (d), we choose $\lambda=1$ , which is located in the region with mobility edges and the number of extended states decreasing with the increase of the quasiperiodic potential strength. We observe a slight difference between distributions at odd and even sites in the long time limit. Finally, in Fig. 1 (e), we have $\lambda=10$, which is in the region with the number of extended states keeping invariant with the increase of $\lambda$. Here, a clear discrepancy between distributions at even and odd sites emerges, maintaining throughout evolution while retaining identical propagation velocities for both sites. Additionally, in Figs. 1(c-e), the wave packet gradually expands over time. After a sufficiently long evolution time, a portion of the wave packets in the center remains localized at their initial positions, while another portion spreads across the entire system, attributed to the presence of localized eigenstates, as shown in Figs. 1(d-e). It is noteworthy that while wave packet expansion occurs in all cases, the time scales of expansion are different.
@@
In order to quantitatively describe the expansion of wave packets and the time scales of the dynamics, we consider the mean-square displacement Abe , Piechon , TongPQ , which is defined as
$\sigma^{2}(t)=\langle\psi(t)|(\hat{X}-j_{c})^{2}|\psi(t)\rangle,$
@@
where $\hat{X}$ is the coordinate operator. In Fig.2, we display the growth of $\sigma^{2}(t)$ under different strengths of quasiperiodic potential. The system with larger quasiperiodic potential strength will expand more slowly: to achieve the same wave packet width, it requires longer time for larger $\lambda$. Although having different time scales, the dynamical behaviors in the large quasiperiodic potential strength region are quite similar. In the inset of Fig.2, we find the lines in the large quasiperiodic potential strength region coincide when we choose a rescaled x-axis ($t/\lambda$). That means the time scales of wave packet dynamics are proportional to the quasiperiodic potential strength $\lambda$. In other words, the quasiperiodic strength can linearly modulate the time scale of wave packet dynamics. Further, we fit the linear interval in Fig.2 and find that $\sigma^{2}(t)\propto t^{2}$ which indicates the transport of wave packet are ballistic all for different quasiperiodic potential strengths.
@@
Figure 2: Growth of $\sigma^{2}(t)$ under different quasiperiodic potential strength. We choose different quasiperiodic potential strength from $\lambda=0.1$ to $\lambda=20$, $L=987$ and $\sigma=1$. We take $100$ quasiperiodic potential realizations in our calculation. In the inset, we have chosen a rescaled x-axis ($t/\lambda$). It is shown that all lines in the large quasiperiodic potential strength region coincide, suggesting that the time scale of the dynamic evolution is proportional to $\lambda$. The dashed line is $\sigma_{2}(t)\propto t^{2}$.
@@ II.3 Particle number distribution
Next we study the time evolution of particle number distribution. By choosing different initial states and counting the number of particles at different sites, we can calculate the two-point correlation function $\hat{\rho}_{ij}=\langle\psi(t)|\hat{c}^{\dagger}_{i}\hat{c}_{j}|\psi(t)\rangle$. The details of the calculation of the two-point correlation function can be found in appendix A. The diagonal part of $\hat{\rho}_{ii}$ gives particle number distribution $\hat{n}_{i}$.
The first quantity we considered is the imbalance ImbalanceA which is defined as:
@@
where $N=N_{\text{even}}+N_{\text{odd}}$ is the total particle number and
$N_{\text{even(odd)}}(t)=\sum_{j=\text{even(odd)}}\langle\psi(t)|\hat{n}_{j}|\psi(t)\rangle$
@@
Figure 3: The evolution of Imbalance $\mathcal{I}(t)$ for different quasiperiodic potential strength. We choose $L=100$ and $N=50$ in our calculation and take $100$ disorder realizations.
@@
In order to detect the effect of the quasiperiodic potential strength on the time scale of the system, we consider another initial state given by
$|\psi(0)\rangle=\prod_{j=1}^{N}\hat{c}^{\dagger}_{j}|0\rangle$ (8)
@@
Figure 4: Time evolution of $N_{h}(t)$ for different disorder strength. We choose $L=100$ and $N=50$ in our calculation and take $100$ disorder realizations.
In the large quasiperiodic potential strength region, the saturation values are the same for different quasiperiodic potential strengths. However, the time scale changes with the increase of the quasiperiodic potential strength. When we rescale the x-axis to $t/\lambda$, we find these curves coincide in large quasiperiodic potential strength region.

### v2

# Types of dynamical behavior in a quasiperiodic mosaic lattice

@@ Abstract
Quasiperiodic mosaic systems with the quasiperiodic potential being added periodically with a fixed lattice interval have attracted significant attention due to their special spectral properties with exactly known mobility edges, which separate localized and extended states. These mobility edges do not vanish even in the region of large quasiperiodic potential strength, although the width of the energy window of extended states decreases with the increase in potential strength and thus becomes very narrow in the limit of strong quasiperiodic disorder. In this paper, we study the dynamics of a quasiperiodic mosaic lattice and unravel its peculiar dynamical properties. By scrutinizing the expansion dynamics of wave packet and the evolution of density distribution, we unveil that the long-time density distribution displays obviously different behaviors at odd and even sites in the region of large quasiperiodic potential strength. Particularly, the time scale of dynamics exhibits an inverse relationship with the quasiperiodic potential strength. To understand these behaviors, we derive an effective Hamiltonian in the large quasiperiodic potential strength region, which is composed of decoupled Hamiltonians defined on the odd and even sites, respectively. While all eigenstates of the effective Hamiltonian defined on even sites are localized, the eigenstates of effective Hamiltonian defined on odd sites include both localized and extended eigenstates. Our results suggest that the effective Hamiltonian can describe the dynamical behaviors well in the large quasiperiodic potential strength region and provides an intuitive framework for understanding the peculiar dynamical behaviors in the quasiperiodic mosaic lattice.
@@ I Introduction
Quasiperiodic systems, as pivotal platforms for investigating Anderson localization Anderson , Anderson50 , Mirlin , Lee , 2D and mobility edges MEAnderson1 , MEAnderson2 , Mott1 , have attracted considerable attention and extensive exploration. Compared to the Anderson model in which the extended-localized transition and mobility edge only exist in three dimension 2D , MEAnderson1 , MEAnderson2 , Mott1 , Experiment1 , Experiment4 , Experiment5 , quasiperiodic systems can host the the extended-localized transition and mobility edge in one dimension AAmodel , AAHmodel , Cai , HYao , MobilityEdge1 , self-consisitent , Experiment2 , Experiment3 , XCXie , XCXie1 , SarmaMobilityEdge1 , SarmaMobilityEdge2 , GaoXL . While the Aubry-Audré (AA) model hosts extended-localized transition AAmodel , AAHmodel at the self-dual point, generalizations of the AA model MobilityEdge1 , self-consisitent , XCXie , XCXie1 , SarmaMobilityEdge1 , SarmaMobilityEdge2 , Deng , longrange , MosaicAA , globaltheory1 , hidden-duality , RG , Critical , DDVu1 , LiuXJ , WangLi or periodically driving the quasiperiodic systems Qin , Driving , Driving1 , Driving2 , Driving3 , Driving4 can give rise to fruitful phenomena of mobility edges. Recently, a class of quasiperiodic mosaic models MosaicAA , in which the quasiperiodic potential periodically occurs with a fixed interval, was proposed as another type of quasiperiodic systems with exact mobility edges. Different from previous models SarmaMobilityEdge1 , SarmaMobilityEdge2 , the quasiperiodic mosaic models have mobility edges even in the large quasiperiodic potential strength region and can host multiple mobility edges. Stimulated by this quasiperiodic mosaic model, many other mosaic systems have been studied and show quite interesting behaviors LiuYX , Longhi , MosaicSystem1 , MosaicSystem2 , MosaicSystem3 , MosaicSystem4 , MosaicSystem5 , Mosaic6 , Dwiputra . Recently, the experimental realization of the quasiperiodic mosaic model was also reported mosaicAAExperiment .
@@
While the diffusion of a wave packet in a uniform lattice is ballistic, the Anderson localization can suppress the diffusion of the wave packet. For a one-dimensional quasiperiodic lattice with mobility edge, for example, the Ganeshan-Pixley-Das Sarma (GPD) model SarmaMobilityEdge2 , there exists an intermediate regime in which the extended and localized states coexist and are separated by the mobility edges. The dynamical behaviors in this region are quite subtle and have drawn considerable attention, showcasing unique phenomena Dynamics1 , Dynamics2 , Dynamics3 . It was shown that the dynamical behaviors of a wave packet in the intermediate regime of the GPD model exhibit a blend of both localized and extended features Dynamics1 , distinguishing them from dynamics in the multifractal region. For quasiperiodic mosaic systems, despite Avila’s global theory GlobalTheory providing an exact expression for the mobility edge, the dynamics of the quasiperiodic mosaic model has not yet been studied and is worth further exploration.
@@
In this paper, we investigate the dynamics of a quasiperiodic mosaic model across a range of quasiperiodic potential strengths and particularly unravel the peculiar dynamical behavior in the large quasiperiodic potential strength limit. Our analysis involves visually tracking the evolution of a Gaussian wave packet and study the evolution of particle number distribution with different initial states. We observe quite distinct behaviors under different quasiperiodic potential strengths. Particularly, in the large quasiperiodic potential strength limit, we observe significant density distribution differences between even and odd sites, with the expansion timescale of the wave packet showing a linear relationship with the quasiperiodic potential strength $\lambda$. To understand the physical original of these peculiar dynamical behaviors, we derive an effective Hamiltonian in the large quasiperiodic potential strength region, indicating that the effective Hamiltonian for the odd and even sites are decoupled. This effective Hamiltonian provides a clear physical picture of the even-odd difference of the dynamical behaviors and effectively captures the scaling behavior of eigenvalues, as well as the property of extended-localized transition. Additionally, analysis of the eigenstates of the effective Hamiltonian yields further insights into the understanding of the spectral structure of the quasiperiodic mosaic model. Our study sheds light on the physical properties of the quasiperiodic mosaic model, unraveling the physical origin of the peculiar dynamical behaviors in the quasiperiodic mosaic model.
@@
The rest of the paper is organized as follows: In Sec.II, we introduce the quasiperiodic mosaic model and its main properties briefly. Then, we study the dynamics of the wave packet expansion and the evolution of particle number distribution with different initial states. In Sec.III, we derive the effective Hamiltonian in the large quasiperiodic potential strength limit and analyze the properties of the effective Hamiltonian, through which we can understand clearly the even-odd difference of the wave packet dynamics. Finally, in Sec. IV, we give a summary of our findings.
@@ II.1 Model
where $\hat{c}_{j}^{\dagger}(\hat{c}_{j})$ is the fermion creation(annihilation) operator and $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the particle number operator at the $j$th site. $\lambda$ is the strength of quasiperiodic potential and $J$ is the hopping amplitude, for which we set $J=1$ as the unit of energy in the following calculation. $\omega$ is an irrational number and we choose $\omega=\frac{\sqrt{5}-1}{2}$. $\phi$ is a random phase which leads to the realization of different quasiperiodic potentials. We note that the quasiperiodic potential is distributed only on even sites.
@@
where $\psi_{n,j}$ represents the $n$th eigenstate’s amplitude on the $j$th site. The fractal dimension $D_{2}$ takes different values at the thermodynamic limit in various regions: while $D_{2}\rightarrow 1$ in the extended region, $D_{2}\rightarrow 0$ in the localized region.
@@
Figure 1: Basic properties and wave packet dynamics of the quasiperiodic mosaic model. (a) The eigenvalue spectrum versus $\lambda$ with the fractal dimension of corresponding eigenstate denoted by $D_{2}$. (b) The proportion of extended states in quasiperiodic mosaic model. (c)-(e) The evolution of wave packet under different quasiperiodic potential strength. From left to right, we choose $\lambda=0.1,1,10$. From top to bottom, we choose $t=0,10,10^{2},10^{3},10^{4}$. The red circle (blue diamond) markers represent the even (odd) wavefunction distribution. We choose the Gaussian wave packet with $\sigma=1$ as the initial state. In (e), the solid(dashed) lines represent the even (odd) wavefunction distribution from the evolution of the effective Hamiltonian. We choose $L=987$ in our calculation and take the average over 100 realizations of the quasiperiodic potential.
@@ II.2 Wave packet dynamics
$|\psi(0)\rangle=C\sum_{j}\exp(-(j-j_{c})^{2}/2\sigma^{2})|j\rangle,$ (6)
@@
To get an intuitive picture of the evolution of wave packets for various $\lambda$, we show the distribution of particle number at different times $N_{j}(t)=\langle\psi(t)|\hat{n}_{j}|\psi(t)\rangle$. The results are shown in Fig.1(c-e), which exhibit the wave packet dynamics across varying degrees of quasiperiodic potential strength. In Fig. 1 (c), with $\lambda=0.1$, all the eigenstates are extended. The wave packet expands and lose all the initial information after a long time evolution. Here we adopted logarithmic coordinates for $N_{j}(t)$. Notably, there are no discernible differences between distributions at even and odd sites. In Fig.1 (d), we choose $\lambda=1$ , which is located in the region with mobility edges and the number of extended states decreasing with the increase of the quasiperiodic potential strength. We observe a slight difference between distributions at odd and even sites in the long time limit. Finally, in Fig. 1(e), we have $\lambda=10$, which is in the region with the number of extended states keeping invariant with the increase of $\lambda$. Here, a clear discrepancy between distributions at even and odd sites emerges, maintaining throughout evolution while retaining identical propagation velocities for both sites. Additionally, in Figs. 1(c)-1(e), the wave packet gradually expands over time. After a sufficiently long evolution time, a portion of the wave packets in the center remains localized at their initial positions, while another portion spreads across the entire system, attributed to the presence of localized eigenstates, as shown in Figs. 1(d)-1(e). It is noteworthy that while wave packet expansion occurs in all cases, the time scales of expansion are different.
@@
To quantitatively describe the expansion of wave packets and the time scales of the dynamics, we consider the mean-square displacement Abe , Piechon , TongPQ , which is defined as
$\sigma^{2}(t)=\langle\psi(t)|(\hat{X}-j_{c})^{2}|\psi(t)\rangle,$ (8)
@@
where $\hat{X}$ is the coordinate operator. In Fig.2, we display the growth of $\sigma^{2}(t)$ under different strengths of quasiperiodic potential. The system with larger quasiperiodic potential strength will expand more slowly: to achieve the same wave packet width, it requires longer time for larger $\lambda$. Although having different time scales, the dynamical behaviors in the large quasiperiodic potential strength region are quite similar. In the inset of Fig.2, we find the lines in the large quasiperiodic potential strength region coincide when we choose a rescaled $x$-axis ($t/\lambda$). That means the time scales of wave packet dynamics are proportional to the quasiperiodic potential strength $\lambda$. In other words, the quasiperiodic strength can linearly modulate the time scale of wave packet dynamics. After a short increase, we find that the linear interval in Fig.2 can be well fitted by $\sigma^{2}(t)\propto t^{2}$ before they reach saturation values, which indicates the transport of the wave packet are ballistic all for different quasiperiodic potential strengths.
@@
Figure 2: Growth of $\sigma^{2}(t)$ under different quasiperiodic potential strength. We choose different quasiperiodic potential strength from $\lambda=0.1$ to $\lambda=20$, $L=987$ and $\sigma=1$. We take the average over 100 realizations of the quasiperiodic potential in our calculation. In the inset, we have chosen a rescaled $x$-axis ($t/\lambda$). It is shown that all lines in the large quasiperiodic potential strength region coincide, suggesting that the time scale of the dynamic evolution is proportional to $\lambda$. The dashed line is $\sigma^{2}(t)\propto t^{2}$.
@@ II.3 Particle number distribution
Next we study the time evolution of particle number distribution for a half-filled system. By choosing different initial states and counting the number of particles at different sites, we can calculate the two-point correlation function $\hat{\rho}_{ij}=\langle\psi(t)|\hat{c}^{\dagger}_{i}\hat{c}_{j}|\psi(t)\rangle$. The details of the calculation of the two-point correlation function can be found in Appendix A. The diagonal part of $\hat{\rho}_{ii}$ gives particle number distribution $\hat{n}_{i}$.
The first quantity we considered is the imbalance ImbalanceA , which is defined as:
@@
where
$N_{\text{even(odd)}}(t)=\sum_{j=\text{even(odd)}}\langle\psi(t)|\hat{n}_{j}|\psi(t)\rangle$ (10)
@@
Figure 3: The evolution of imbalance $\mathcal{I}(t)$ for different quasiperiodic potential strength. We choose $L=100$ in our calculation and take the average over 100 realizations of the quasiperiodic potential. Besides, we place different markers on the lines with different $\lambda$ to differentiate them.
@@
To detect the effect of the quasiperiodic potential strength on the time scale of the system, we consider another initial state given by
$|\psi(0)\rangle=\prod_{j=1}^{L/2}\hat{c}^{\dagger}_{j}|0\rangle$ (12)
@@
Figure 4: Time evolution of $N_{h}(t)$ for different disorder strength. We choose $L=100$ in our calculation and take the average over 100 realizations of the quasiperiodic potential.
In the large quasiperiodic potential strength region, the saturation values are the same for different quasiperiodic potential strengths. This is because the underlying structures (fractions of localized or extended eigenstates) remain the same and independent of $\lambda$ in this region. However, the time scale changes with the increase of the quasiperiodic potential strength. When we rescale the x-axis to $t/\lambda$, we find these curves coincide in the large quasiperiodic potential strength region.